OpenStudy (automaticloveletter):

Helene paid back $100 in Month 1 of her loan. In each month after that, Helene paid back $50. Write an explicit formula and a recursive formula that shows f(n), the total amount Helene had paid back by Month n.

OpenStudy (amistre64):

what is your concept of the forms for these equations?

OpenStudy (automaticloveletter):

what do you mean? recursive and explicit?

OpenStudy (amistre64):

yes

OpenStudy (automaticloveletter):

recursive is f(x)= (n-1) something like that ?

OpenStudy (amistre64):

recurse defines the current amount by adding to the last months amount;and it has to define a starting month value ... (starting month) = K (this month) = (last month) + (some payment) ------------- M_{1} = 100 M_{n} = M{n-1}+ 50

OpenStudy (amistre64):

f(n) and f(n-1) i spose are the proper notation for this question

OpenStudy (amistre64):

can you fix the notation?

OpenStudy (automaticloveletter):

so f(n)= f(n-1)+50?

OpenStudy (amistre64):

and the starting value?

OpenStudy (automaticloveletter):

f(1)=100?

OpenStudy (amistre64):

good, then thats our setup defined recursivly f(n) = f(n-1) + 50, f(1) = 100 now we could easily start at 100, and add 50 a month for however many months ... right?

OpenStudy (automaticloveletter):

yes

OpenStudy (amistre64):

out explicit formula simply does that, it starts at 100, and we add on 50 for however many months we pay ... f(n) = 100 + 50(number of months) now the number of months here has a little oddity to it, simply a correction factor. f(1) = 100, only of 50*(number of months) is 0 if n represents the number of months, then how do we make that go to zero when n=1?

OpenStudy (automaticloveletter):

im confused, do we add 100 & 50 to get f(1)'s answer?

OpenStudy (amistre64):

hmm, im not sure how your material wants you to define the explicit equation, but the 2 methods i have in my head are mathically equivalent. we are adding up a bunch of 50s, a 50 for each month 50n is therefore the amount we have paid over n months, does this make sense? in 3 months we have paid out at least 50(3)=150 dollars

OpenStudy (automaticloveletter):

yes that makes sense

OpenStudy (amistre64):

then f(n) = 50n is almost right for us ... except for 1 case. f(1) = 50(1) is not equal to 100. how much are we off by? how do we correct this?

OpenStudy (automaticloveletter):

50?

OpenStudy (amistre64):

yeah, so lets add 50 f(n)= 50 + 50n would satisfy the results. another form to write it in is: f(n)= 100 + 50(n-1) or f(n)=50(n+1) all of these are just algebraic manipulations of the same equation, they are all equivalent

OpenStudy (automaticloveletter):

can you help me with some more?

OpenStudy (amistre64):

maybe

OpenStudy (automaticloveletter):

so it basically says let f(x)=2x-1 g(x)=3x &h(x)=x^2+1

OpenStudy (automaticloveletter):

and the first question says f(g(-3))

OpenStudy (amistre64):

how would you determine g(-3) ?

OpenStudy (automaticloveletter):

g(x)=3x?

OpenStudy (amistre64):

g(x)=3x g(-3)....notice that x is replaced by -3 as the input value; so what is the value of g, when x=-3?

OpenStudy (automaticloveletter):

uhh 3x?

OpenStudy (amistre64):

3x is not replacing x by -3 .... you have to know how to work a function.

OpenStudy (amistre64):

when x=-3, what value do we get for the function of g? \[\large g(\underbrace{\color{red}{x}}_{-3} ) = 3\underbrace{\color{red}{x}}_{-3}\]

OpenStudy (automaticloveletter):

No clue... -3?

OpenStudy (amistre64):

then how do we read 3x? what does 3x mean?

OpenStudy (automaticloveletter):

g(-3)*3x ?

OpenStudy (automaticloveletter):

I don't know I'm sorry...

OpenStudy (amistre64):

g is the name of a function, its value depends on the variable x: g(x) reads:g of x 3x defines how g and x are related. 3x reads: 3, times x: x+x+x if x=-3, then 3(-3) = -3-3-3, what is 3, times -3?

OpenStudy (automaticloveletter):

-9.

OpenStudy (amistre64):

then g(-3)= -9 now the bummer of this is that we have to use the same logic to define f(-9)

OpenStudy (amistre64):

the function f,defined by x: f(x) is related by the expression: 2x-1 if x is equal to -9, then what is the value of f(-9) ?

OpenStudy (amistre64):

\[f(\color{red}{g(-3)})\implies f(\color{red}{-9})\implies 2(\color{red}{-9})-1\]

OpenStudy (automaticloveletter):

you mean 2*-9-1.. -19?

OpenStudy (amistre64):

yep

OpenStudy (automaticloveletter):

so it's -19?

OpenStudy (amistre64):

that was what we determined it to be, yes